OpenStudy (anonymous):

Find dy/dx by implicit differentiation. x^2 + xy + y^2 = 4. Okay, so I know that I have to differentiate the problem so that x^2 becomes 2x and y^2 becomes 2y • dy/dx and 4 becomes 0. What should I do with xy? I know I need to use the product rule, but my answer just isn't matching up to the answer I was given.

OpenStudy (bibby):

d/dx uv = u'v + uv' where v' would be y' or dy/dx

OpenStudy (anonymous):

Okay, so it would be x'y + xy' which would then equal y + x, right?

OpenStudy (bibby):

y+xy'

OpenStudy (anonymous):

Okay, so it would be x + y times dy/dx, right?

OpenStudy (bibby):

how'd you get that?

OpenStudy (anonymous):

Because y' = dy/dx and the problem wants us to work with dy/dx.

OpenStudy (solomonzelman):

this is what is happening, you are differentiating term by term (each term separately). You are differentiation X's and Y's all the same y, BUT whenever you are differentiating Y you are multiplying that times dy/dx (or times y', -- whatever notation satisfies you) as you differentiate all terms, solve for y' (or for dy/dx -- whatever you wrong the derivative as)

OpenStudy (bibby):

my point is \(\dfrac{d}{dx}xy=x'y+xy'=y+x\dfrac{dy}{dx}\) you flipped the x and y

OpenStudy (anonymous):

Ohhh, oops! I didn't mean to do that, I'm sorry!

OpenStudy (bibby):

don't be sorry, just don't do it again

OpenStudy (bibby):

jk

OpenStudy (solomonzelman):

"wrong" in last parenthesis shouldn't be there "it is supposed to be "write"

OpenStudy (anonymous):

@SolomonZelman Wait, what do you mean by your last statement?

OpenStudy (solomonzelman):

I am correcting my first one

OpenStudy (bibby):

he typed wrong instead of wrote

OpenStudy (anonymous):

Ohhh, okay

OpenStudy (solomonzelman):

yeah

OpenStudy (anonymous):

Okay, so I'm at the point where I have this: 2x + (y + x•dy/dx) + 2y = 0. Is this correct so far?

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

the derivative of y^2 is also multiplied times y'

OpenStudy (anonymous):

Dang it, you're right. So it would be 2y times dy/dx?

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

Okay, so I have 2x + (y + x•dy/dx) + 2y•dy/dx= 0. I know I need to get dy/dx alone on one side. I can subtract 2x at least, but what do I do about the rest?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2x+y+\frac{dy}{dx}x + \frac{dy}{dx}2 y = 0 }\) (you can use my code, ``` \(\large\color{black}{ \displaystyle 2x+y+\frac{dy}{dx}x + \frac{dy}{dx}2 y = 0 }\) ``` )

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

now isolate \(\large\color{black}{ \displaystyle \frac{dy}{dx} }\)

OpenStudy (anonymous):

Okay, so I've got this so far: x•dy/dx + 2y•dy/dx = =2x - y. I can divide to get 2y and x on the other side. I'm a bit confused at this part.

OpenStudy (solomonzelman):

-2x ? on the right

OpenStudy (solomonzelman):

you are subtracting 2x from both sides aren't you?

OpenStudy (anonymous):

-2x, my bad!

OpenStudy (solomonzelman):

yes, then factor out of dy/dx on the left side

OpenStudy (anonymous):

That was a typo on my part.

OpenStudy (solomonzelman):

it's fine

OpenStudy (anonymous):

Okay, but if I divide by both x and 2y to isolate dy/dx, won't it become dy/dx + dy/dx? Or am I missing something?

OpenStudy (solomonzelman):

I would say messing something, because I don't get what you mean... \(\large\color{black}{ \displaystyle 2x+y+\frac{dy}{dx}x + \frac{dy}{dx}2 y = 0 }\) \(\large\color{black}{ \displaystyle \frac{dy}{dx}x + \frac{dy}{dx}2 y = -2x-y }\) \(\large\color{black}{ \displaystyle \frac{dy}{dx}(x + 2 y )= -2x-y }\) and then the last step

OpenStudy (anonymous):

Ooooohhhhhh! I get it now!

OpenStudy (anonymous):

So then it would become dy/dx = -2x - y / x + 2y!

OpenStudy (solomonzelman):

yes, but don't put the exclamation mark

OpenStudy (solomonzelman):

and put the parenthesis

OpenStudy (anonymous):

Aha, sorry about that. However I do have one more question. The answer for this problem that I was given is (2x + y)/(2y - x). How come these answers don't match up?

OpenStudy (solomonzelman):

it is wrong

OpenStudy (anonymous):

My answer or the answer I was given?

OpenStudy (solomonzelman):

we did all steps together. Our answer is fine

OpenStudy (solomonzelman):

the answer you were given intially is wrong

OpenStudy (anonymous):

Ahhh, okay. Thank you all who helped me so much, I really appreciate it.

OpenStudy (solomonzelman):

Confirmation http://www.wolframalpha.com/input/?i=derivative+x%5E2+%2B+xy+%2B+y%5E2+%3D+4.

OpenStudy (solomonzelman):

yes, our answer is right. Wolfram also agrees to us.

OpenStudy (anonymous):

Okay. Thank you all again so much!

OpenStudy (solomonzelman):

Sure!